OpenStudy (anonymous):

Help for a metal?!?! The art museum had a total of 224 visitors on Tuesday. Visitors older than 18 paid $12 for admission. Visitors 18 years or younger paid $8 for admission. The museum collected a total of $2,520. How many visitors were 18 or younger?

OpenStudy (anonymous):

I don't like metal ;c

OpenStudy (laughoutloud):

42

OpenStudy (cj49):

let visitors older den 18 be x and younger than 18 b y 12x+8y=2520 simplyfy the equation. den substitute the values.

OpenStudy (anonymous):

42 people were under 18,

OpenStudy (anonymous):

224 total people. x<12 y> 8 2520 total made. 12x + 8y = 2520

OpenStudy (anonymous):

laughoutloud can you explain how it's 42?

OpenStudy (anonymous):

8u + 12o = 2520 u + o = 224 - o = u - 224 8u + 12(u - 224) = 2520 8u + 12u - 2688 = 2520 20u - 2688 = 2520 20u = 5208 u = 260.4

OpenStudy (anonymous):

Same with lupita

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (anonymous):

yw welcome

OpenStudy (anonymous):

Can one of you help me with another question?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Thank you! Here is the next question: Solve the following system of equations: -2x + y = 1 -4x + y = -1

OpenStudy (anonymous):

@lupita1995

OpenStudy (anonymous):

@lupita1995

OpenStudy (anonymous):

You there?